Whitman Company has just completed its first year of operations. The company’s absorption costing income statement for the year follows:

Whitman Company
Income Statement
Sales (38,000 units × $43.60 per unit) $ 1,656,800
Cost of goods sold (38,000 units × $22 per unit) 836,000
Gross margin 820,800
Selling and administrative expenses 437,000
Net operating income $ 383,800

The company’s selling and administrative expenses consist of $285,000 per year in fixed expenses and $4 per unit sold in variable expenses. The $22 unit product cost given above is computed as follows:

Direct materials $ 10
Direct labor 5
Variable manufacturing overhead 3
Fixed manufacturing overhead ($204,000 ÷ 51,000 units) 4
Absorption costing unit product cost $ 22

Required:

1. Redo the company’s income statement in the contribution format using variable costing.

2. Reconcile any difference between the net operating income on your variable costing income statement and the net operating income on the absorption costing income statement above.
Answer #1
1
Whitman Company
Variable Costing Income Statement
Sales 1656800
Variable expenses:
Variable cost of goods sold 684000 =38000*(10+5+3)
Variable selling and administrative expenses 152000 =38000*4
836000
Contribution margin 820800
Fixed expenses:
Fixed manufacturing overhead 204000
Fixed selling and administrative expenses 285000
489000
Net operating income 331800
2
Variable costing net operating income 331800
Add: Fixed manufacturing overhead deferred in
inventory under absorption costing
52000 =(51000-38000)*4
Absorption costing net operating income 383800